Problem 5-1 Simple Interest versus Compound Interest (LO1) 

First City Bank pays 7 percent simple interest on its savings account balances, whereas Second City Bank pays 7 percent interest compounded annually. 


If you made a $56,000 deposit in each bank, how much more money would you earn from your Second City Bank account at the end of 9 years? (Do not round Intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) 

Difference in accounts =  $ _______ 
Answer #1

Calculate the difference amount as follows:

B 362 Particulars Amount 363 First bank $91,280.00 364 second bank $102,953.72 365 Difference $11,673.72

Formulas:

А B 362 Particulars Amount 363 First bank =56000+(56000*7%*9) 364 second bank =FV(7%,9,-56000) 365 Difference =B364-B363